OpenStudy (anonymous):

The length of a rectangle is 6 m less than twice its width. The area of the rectangle is 140 m^2. What is the length of the shorter side of the rectangle?

OpenStudy (anonymous):

The width = x The length - 2x - 6 Area = length * width 140 = x(2x-6) 140 = 2x^2 - 6x 0 = 2x^2 - 6x - 140 0 = 2(x^2 - 3x - 70) 0 = 2(x - 10)(x + 7) x = 10 or -7 Width = 10 Length = 2(10) - 6 = 14 The shortest side is the width which is 10